OpenStudy (anonymous):

simplify radical expression square root of 15 divided by 5 square root of 20 please show work.

OpenStudy (yanasidlinskiy):

Can you draw it out??

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ 15\implies 3\cdot 5\qquad 20\implies 2\cdot 2\cdot 5}}\implies 2^2\cdot 5\\ \quad \\ \quad \\ \cfrac{\sqrt{15}}{5\sqrt{20}}\implies \cfrac{\sqrt{3\cdot 5}}{5\sqrt{2^2\cdot 5}}\implies \cfrac{\sqrt{3}\cdot \cancel{\sqrt{5}}}{5\sqrt{2^2}\cdot \cancel{\sqrt{5}}}\)

OpenStudy (jdoe0001):

?

OpenStudy (anonymous):

it says the answer comes out to be square root of 15 divided by 3. but i dont know how it comes up with that answer.

OpenStudy (anonymous):

ok i got it i was looking at the wrong problem thank you so much (:

OpenStudy (jdoe0001):

yw